# A Bayesian latent Gaussian conditional autoregressive copula model for analyzing spatially-varying trends in rainfall

### v1

@@ 4.1 CAR copula
Let $\tilde{\bm{X}}=(\tilde{X}_{1},\ldots,\tilde{X}_{n})^{\prime}$ be a random vector following a CAR model defined in Section 2.2, where the adjacency matrix is constructed using the neighborhood structure of the meteorological subdivisions of mainland India. Given the spatial autocorrelation parameter $\rho$ and fixed unit marginal variance ($\sigma^{2}=1$), the joint distribution of $\tilde{\bm{X}}$ is given by $\tilde{\bm{X}}\sim\textrm{Normal}_{n}(\bm{0},[\bm{M}-\rho\bm{W}]^{-1})$, where $\bm{M}$ and $\bm{W}$ are as defined in Section 2.2 and $\textrm{Normal}_{n}$ denotes an $n$-variate normal distribution. Suppose the diagonal elements of the matrix $[\bm{M}-\rho\bm{W}]^{-1}$ are given by $d_{i}$ for $i=1,\ldots,n$ and $\bm{\Delta}=\textrm{diag}(d_{1},\ldots,d_{n})$. Then, $\bm{\Delta}^{-1/2}\tilde{\bm{X}}\sim\textrm{Normal}_{n}(\bm{0},\bm{\Delta}^{-1/2}[\bm{M}-\rho\bm{W}]^{-1}\bm{\Delta}^{-1/2})$ and its corresponding marginal distributions are $\textrm{Normal}(0,1)$. Denoting $U_{i}=\Phi(d_{i}^{-1/2}\tilde{X}_{i})$ for $i=1,\ldots,n$, where $\Phi(\cdot)$ denotes the standard normal CDF, we call the joint distribution of $\bm{U}=(U_{1},\ldots,U_{n})^{\prime}$ a CAR copula with adjacency matrix $\bm{W}$ and spatial autocorrelation parameter $\rho$. The joint CDF of $\bm{U}$ is given by
@@ 4.2 Data layer modeling
We recall the data setup from Section 3, where the annual total rainfall data are denoted by $Y_{it}$ for $i=1,\ldots,n=34$ meteorological subdivisions of India and for $t=1,\ldots,T=64$ years between 1951 and 2014.
@@
where $t^{*}=(t-m_{t})/s_{t}$ with $m_{t}=T^{-1}\sum_{t=1}^{T}t$ and $s_{t}^{2}=T^{-1}\sum_{t=1}^{T}t^{2}-m_{t}^{2}$. Henceforth, we denote the CDF of the gamma distribution in (11) as $F(\cdot;a_{i},b_{i},c_{i})$. Thus, $U_{it}=F(Y_{it};a_{i},b_{i},c_{i})\sim\textrm{Uniform}(0,1)$. Let the vectors of $a_{i}$s, $b_{i}$s, and $c_{i}$s be denoted by $\bm{a}$, $\bm{b}$, and $\bm{c}$, respectively, as described in Section 3. Conditioning on $\bm{a}$, $\bm{b}$, and $\bm{c}$, while we assume independence across years, we model the spatial dependence of $\bm{U}_{t}=(U_{1t},\ldots,U_{nt})^{\prime}$ in a copula framework as described in Section 4.1. Overall, the density of $\bm{Y}_{t}=(Y_{1t},\ldots,Y_{nt})$ given $\bm{a}$, $\bm{b}$, $\bm{c}$, and $\rho$ is
@@
×$\times\prod_{i=1}^{n}f(y_{it};a_{i},b_{i},c_{i})\big{/}\prod_{i=1}^{n}\phi(\Phi^{-1}(u_{it}))~{}~{},$
where $u_{it}=F(y_{it};a_{i},b_{i},c_{i})$, $f(y_{it};a_{i},b_{i},c_{i})$ denotes the density corresponding to the gamma distribution in (11). Further, assuming independence across time, the joint density of the full data is $f_{n}(\bm{y}_{1},\ldots,\bm{y}_{T}|\bm{a},\bm{b},\bm{c},\rho)=\prod_{t=1}^{T}f_{n}(\bm{y}_{t}|\bm{a},\bm{b},\bm{c},\rho)$.
@@ 4.3 Bayesian latent Gaussian model
In this subsection, we subsequently model the areally-varying coefficients $a_{i}$s, $b_{i}$s, and $c_{i}$s. The range of these coefficients are $a_{i}>0,b_{i}>0,c_{i}\in\mathbb{R}$. We thus consider the logarithmic transformation of the first two and define $a_{i}^{*}=\log(a_{i})$ and $b_{i}^{*}=\log(b_{i})$ and choose different dependent or independent Gaussian priors for them as follows.
Case 1. We assume independent and identically distributed (iid) Gaussian priors to the parameters as follows
@@
for all $i=1,2,\dots,n$. Here, $\mu_{a}$ and $\sigma_{a}^{2}$ represent the overall average of $a_{i}^{*}$s and spatial variability of $a_{i}^{*}$s considering mainland India. Similarly, $\mu_{b}$ and $\sigma_{b}^{2}$ represent the overall average of $b_{i}^{*}$s and spatial variability of $b_{i}^{*}$s, and $\mu_{c}$ and $\sigma_{c}^{2}$ represent the overall average of $c_{i}$s and spatial variability of $c_{i}$s. The prior choice in (13) does not allow a spatial dependence structure visible from Figure 2. Besides, the rainfall patterns of two neighboring regions usually behave similarly for natural climatic reasons. Hence, it is likely that (13) would underperform compared to a model that appropriately considers a spatial dependence structure, and the spatial smoothing of the parameters is controlled by the data only.
Case 2. Considering the potential geographical relationships of the rainfall pattern for two nearby meteorological subdivisions, we consider CAR prior specifications for the parameters as follows
@@
Case 3. While the spatial correlation structure in (14) is more general, estimating the hyperparameters $\rho_{a},\rho_{b},$ and $\rho_{c}$ is usually challenging due to high posterior uncertainty. Thus, as a special case, we consider Intrinsic CAR (ICAR) [58] prior specifications for the parameters as follows
$\begin{split}\bm{a}^{*}=(\log(a_{1}),,\dots,\log(a_{n}))^{\top}&{\sim}\text{Normal}_{n}(\mu_{a}\bm{1},\sigma_{a}^{2}[\bm{M}-\bm{W}]^{-1}),\\ \bm{b}^{*}=(\log(b_{1}),,\dots,\log(b_{n}))^{\top}&{\sim}\text{Normal}_{n}(\mu_{b}\bm{1},\sigma_{b}^{2}[\bm{M}-\bm{W}]^{-1}),\\ \bm{c}=(c_{1},\dots,c_{n})^{\top}&{\sim}\text{Normal}_{n}(\mu_{c}\bm{1},\sigma_{c}^{2}[\bm{M}-\bm{W}]^{-1}),\end{split}$ (15)
@@
Finally, we choose a weakly-informative $\textrm{Uniform}(0,1)$ prior for $\rho$. For the hyperparameters, we again choose weakly-informative hyperpriors
@@
Altogether, (12), (14), and (16) specify a fully Bayesian hierarchical model, and we henceforth call it the CAR-CAR model due to a CAR structure at the both data and prior layers. Similarly, we call the model specified by (12), (15), and (16) (ignoring the hyperpriors for $\rho_{a}$, $\rho_{b}$, and $\rho_{b}$) as CAR-ICAR model, and the model specified by (12), (15), and (16) as CAR-Indep model.
@@ 5 Bayesian computation
This section discusses the MCMC sampling steps for the CAR-CAR model. The sampling for the other models mentioned in the previous section can be done simply by ignoring some update steps. For example, the computation for the CAR-ICAR model can be done just by ignoring the updates of $\rho_{a},\rho_{b},\text{and}~{}\rho_{c}$ and setting them to one.
@@ 5.1 Gibbs sampling steps for the CAR-CAR model
The set of parameters and hyperparameters in the model are
@@
which includes some parameter vectors, and the rest are single parameters or hyperparameters. We update each of the vectors $\bm{a}^{*}$, $\bm{b}^{*}$, and $\bm{c}$ simultaneously using a block MH algorithm. Here, by $Y_{it}^{*}$s, we denote the missing data that needs to be imputed within the MCMC steps. The update steps within a Gibbs sampler are as follows.
  • Updating $\bm{a}^{*}$: Sample $\bm{a}^{*}$ via Metropolis-Hasting algorithm from
@@
  • Updating $\bm{b}^{*}$: Similar to $\bm{a}^{*}$, sample $\bm{b}^{*}$ jointly via Metropolis-Hasting algorithm from $f(\bm{b}^{*}|\bm{Y}_{1},\ldots,\bm{Y}_{T},\bm{a}^{*},\bm{c},\rho,\mu_{b},\sigma_{b}^{2},\rho_{b})$.
  • Updating $\bm{c}$: Similar to $\bm{a}^{*}$, and $\bm{b}^{*}$, sample $\bm{c}$ jointly via Metropolis-Hasting algorithm from $f(\bm{c}|\bm{Y}_{1},\ldots,\bm{Y}_{T},\bm{a}^{*},\bm{b}^{*},\rho,\mu_{c},\sigma_{c}^{2},\rho_{c})$
  • Updating $\rho$: Sample $\rho$ via the Metropolis-Hasting algorithm from
@@
  • Updating $\rho_{a}$: Sample $\rho_{a}$ via the Metropolis-Hasting algorithm from
@@
  • Updating $\rho_{b}$: Similarly, sample $\rho_{b}$ from $f(\rho_{b}|\bm{b}^{*},\mu_{b},\sigma_{b}^{2})$ via Metropolis-Hasting algorithm as $\rho_{a}$.
  • Updating $\rho_{c}$: Similarly, sample $\rho_{b}$ from $f(\rho_{c}|\bm{c},\mu_{c},\sigma_{c}^{2})$ via Metropolis-Hasting algorithm as $\rho_{a}$ and $\rho_{b}$.
  • Updating $Y^{*}_{it}s$: For a time point $t$, if $\mathcal{I}$ denotes the region indexes where the data are missing, we construct $\bm{Y}_{t}^{*}$ (set of missing values) based on $i\in\mathcal{I}$. We construct the vector for the rest of the indexes as $\widetilde{\bm{Y}}_{t}$, i.e., based on all $i\notin\mathcal{I}$. Further, we transform all the components to $\textrm{Normal}(0,1)$ scale by the $\Phi^{-1}(F(Y_{it};a_{i},b_{i},c_{i}))$ transformation. Given that the transformed data follows a multivariate normal distribution, we can simulate the observations corresponding to the missing locations using the standard result of the conditional distribution for multivariate normal vectors and then transform back to the original scale using the Gaussian CDF and the gamma quantile function.

### v2

@@ 4.1 CAR copula
Let $\tilde{\bm{X}}=(\tilde{X}_{1},\ldots,\tilde{X}_{n})^{\prime}$ be a random vector following a CAR model defined in Section 2.2, where the adjacency matrix is constructed using the neighborhood structure of the meteorological subdivisions of mainland India. Given the spatial autocorrelation parameter $\rho$ and fixed unit marginal variance ($\sigma^{2}=1$), the joint distribution of $\tilde{\bm{X}}$ is given by $\tilde{\bm{X}}\sim\textrm{Normal}_{n}(\bm{0},[\bm{M}-\rho\bm{W}]^{-1})$, where $\bm{M}$ and $\bm{W}$ are as defined in Section 2.2 and $\textrm{Normal}_{n}$ denotes an $n$-variate normal distribution. Suppose the diagonal elements of the matrix $\bm{M}-\rho\bm{W}]^{-1}$ are given by $d_{i}$ for $i=1,\ldots,n$ and $\bm{\Delta}=\textrm{diag}(d_{1},\ldots,d_{n})$. Then, $\bm{\Delta}^{-1/2}\tilde{\bm{X}}\sim\textrm{Normal}_{n}(\bm{0},\bm{\Delta}^{-1/2}[\bm{M}-\rho\bm{W}]^{-1}\bm{\Delta}^{-1/2})$ and its corresponding marginal distributions are $\textrm{Normal}(0,1)$. Denoting $U_{i}=\Phi(d_{i}^{-1/2}\tilde{X}_{i})$ for $i=1,\ldots,n$, where $\Phi(\cdot)$ denotes the standard normal CDF, we call the joint distribution of $\bm{U}=(U_{1},\ldots,U_{n})^{\prime}$ a CAR copula with adjacency matrix $\bm{W}$ and spatial autocorrelation parameter $\rho$. The joint CDF of $\bm{U}$ is given by
@@
In our implementation, the inverse of $\bm{M}-\rho\bm{W}$ is not explicitly required, except to obtain $\bm{\Delta}$ required for scaling the copula. Rather than computing the full covariance matrix, the diagonals can be computed efficiently from the Cholesky factorization of $\bm{M}-\rho\bm{W}$. Specifically, if $\bm{M}-\rho\bm{W}=\bm{LL}^{\intercal}$, solving $\bm{LZ}=\bm{I}$ yields $\bm{Z}=\bm{L}^{-1}$. The diagonal matrix is further obtained as $\bm{\Delta}=\textrm{diag}(\sum_{i}Z_{i1}^{2},\ldots,\sum_{i}Z_{in}^{2})$. This avoids the $\mathcal{O}(n^{3})$ cost of a full inversion, requiring only $\mathcal{O}(n^{2})$ operations.
@@ 4.2 Data layer modeling
We recall the data setup from Section 3, where the monsoon total rainfall data are denoted by $Y_{it}$ for $i=1,\ldots,n=34$ meteorological subdivisions of India and for $t=1,\ldots,T=64$ years between 1951 and 2014.
@@
where $t^{*}=(t-m_{t})/s_{t}$ with $m_{t}=T^{-1}\sum_{t=1}^{T}t$ and $s_{t}^{2}=T^{-1}\sum_{t=1}^{T}t^{2}-m_{t}^{2}$. Henceforth, we denote the CDF of the gamma distribution in (12) as $F(\cdot;a_{i},b_{i},c_{i})$. Thus, $U_{it}=F(Y_{it};a_{i},b_{i},c_{i})\sim\textrm{Uniform}(0,1)$. Let the vectors of $a_{i}$s, $b_{i}$s, and $c_{i}$s be denoted by $\bm{a}$, $\bm{b}$, and $\bm{c}$, respectively, as described in Section 3. Conditioning on $\bm{a}$, $\bm{b}$, and $\bm{c}$, while we assume independence across years, we model the spatial dependence of $\bm{U}_{t}=(U_{1t},\ldots,U_{nt})^{\intercal}$ in a copula framework as described in Section 4.1. Overall, the density of $\bm{Y}_{t}=(Y_{1t},\ldots,Y_{nt})^{\intercal}$ given $\bm{a}$, $\bm{b}$, $\bm{c}$, and $\rho$ is
@@
$\times\prod_{i=1}^{n}f(y_{it};a_{i},b_{i},c_{i})\big/\prod_{i=1}^{n}\phi(\Phi^{-1}(u_{it}))~~,$
where $u_{it}=F(y_{it};a_{i},b_{i},c_{i})$, $f(y_{it};a_{i},b_{i},c_{i})$ denotes the density corresponding to the gamma distribution in (12). Further, assuming independence across time, the joint density of the full data is $f_{n}(\bm{y}_{1},\ldots,\bm{y}_{T}|\bm{a},\bm{b},\bm{c},\rho)=\prod_{t=1}^{T}f_{n}(\bm{y}_{t}|\bm{a},\bm{b},\bm{c},\rho)$. To assess whether the spatial dependence structure is stationary over time (i.e., whether the dependence parameter $\rho$ varies systematically across years), we compute year-specific, smoothed maximum-likelihood estimates of $\rho$ using a moving-average procedure. Smoothed maximum likelihood estimates of $\rho$ remain consistently high with no systematic temporal trend, supporting our choice of the same $\rho$ across years. The details and figure are provided in Section 2 of the Supplementary Material.
In this proposed Bayesian framework, missing rainfall observations are treated as additional unknown quantities and imputed within the MCMC sampling scheme, conditional on the observed data and model parameters. This procedure yields valid posterior inference under the assumption that the missingness mechanism is ignorable [Rubin1976, little2019statistical], which holds when data are either Missing Completely at Random (MCAR) or Missing At Random (MAR). We do not model the missingness mechanism explicitly, and thus do not accommodate Not Missing At Random (NMAR) scenarios, where the probability of missingness depends on unobserved values. In practice, MAR is the most plausible assumption for our dataset, as missing entries typically arise from administrative or instrumental recording failures rather than systematic dependence on unobserved rainfall values. We adopt the partitioning approach of [rue2005gaussian] for handling missing data, which is more efficient than directly inverting the precision matrix $\bm{Q}$. The computational cost of our rainfall application with $n=34$ regions is negligible, but these techniques ensure scalability to larger spatial dimensions.
@@ 4.3 Bayesian latent Gaussian model
In this subsection, we subsequently model the SVCs $a_{i}$s, $b_{i}$s, and $c_{i}$s. The range of these coefficients are $a_{i}>0,b_{i}>0,c_{i}\in\mathbb{R}$. We thus consider the logarithmic transformation of the first two and define $a_{i}^{*}=\log(a_{i})$ and $b_{i}^{*}=\log(b_{i})$ and choose different dependent or independent Gaussian priors for them as follows.
Independent priors (Indep): We assume independent and identically distributed (iid) Gaussian priors to the parameters as follows
@@
for all $i=1,2,\dots,n$. Here, $\mu_{a}$ and $\sigma_{a}^{2}$ represent the overall average of $a_{i}^{*}$s and the spatial variability of $a_{i}^{*}$s considering mainland India. Similarly, $\mu_{b}$ and $\sigma_{b}^{2}$ represent the overall average of $b_{i}^{*}$s and spatial variability of $b_{i}^{*}$s, and $\mu_{c}$ and $\sigma_{c}^{2}$ represent the overall average of $c_{i}$s and spatial variability of $c_{i}$s. The prior choice in (14) does not allow a spatial dependence structure visible from Figure 2. Besides, the rainfall patterns of two neighboring regions usually behave similarly for natural climatic reasons. Hence, it is likely that (14) would underperform compared to a model that appropriately considers a spatial dependence structure, and the spatial smoothing of the parameters is controlled by the data only.
CAR priors (CAR): Considering the potential geographical relationships of the rainfall pattern for two nearby meteorological subdivisions, we consider CAR prior specifications for the parameters as follows
@@
Intrinsic CAR priors (ICAR): While the spatial correlation structure in (15) is more general, estimating the hyperparameters $\rho_{a},\rho_{b},$ and $\rho_{c}$ is usually challenging due to high posterior uncertainty. Thus, as a special case, we consider Intrinsic CAR (ICAR) [freni2018note] prior specifications for the parameters as follows
$\begin{split}\bm{a}^{*}=(\log(a_{1}),,\dots,\log(a_{n}))^{\intercal}&{\sim}\text{Normal}_{n}(\mu_{a}\bm{1},\sigma_{a}^{2}[\bm{M}-\bm{W}]^{-1}),\\ \bm{b}^{*}=(\log(b_{1}),,\dots,\log(b_{n}))^{\intercal}&{\sim}\text{Normal}_{n}(\mu_{b}\bm{1},\sigma_{b}^{2}[\bm{M}-\bm{W}]^{-1}),\\ \bm{c}=(c_{1},\dots,c_{n})^{\intercal}&{\sim}\text{Normal}_{n}(\mu_{c}\bm{1},\sigma_{c}^{2}[\bm{M}-\bm{W}]^{-1}),\end{split}$ (16)
@@
Finally, we choose a weakly-informative $\textrm{Uniform}(0,1)$ prior for $\rho$. For the hyperparameters, we again choose weakly informative hyperpriors
@@
Altogether, (13), (15), and (17) specify a fully Bayesian hierarchical model, and we henceforth call it the CAR-CAR model due to a CAR structure at both data and prior layers. Similarly, we call the model specified by (13), (16), and (17) (ignoring the hyperpriors for $\rho_{a}$, $\rho_{b}$, and $\rho_{b}$) as CAR-ICAR model, and the model specified by (13), (14), and (17) as CAR-Indep model. For model comparison purposes, we consider several competing models by setting $\rho=0$, i.e., the independence at the data layer, and combining with different cases of prior specifications. We refer to them as Indep-Indep, Indep-CAR, and Indep-ICAR models, respectively.
@@ 5 Bayesian computation
This section discusses the MCMC sampling steps for the CAR-CAR model. Sampling for the other models mentioned in the previous section can be performed simply by ignoring specific update steps. For example, the computation for the CAR-ICAR model can be done just by ignoring the updates of $\rho_{a},\rho_{b},\text{and}~\rho_{c}$ and setting them to one.
@@ 5.1 Gibbs sampling steps for the CAR-CAR model
The set of parameters and hyperparameters in the model is
@@
which includes parameter vectors, single parameters, and hyperparameters. We update each of the vectors $\bm{a}^{*}$, $\bm{b}^{*}$, and $\bm{c}$ simultaneously using a block MH algorithm. Here, by $Y_{it}^{*}$s, we denote the missing data that needs to be imputed within the MCMC steps. The update steps within a Gibbs sampler are as follows.
  • Updating $\bm{a}^{*}$: Sample $\bm{a}^{*}$ via Metropolis-Hastings algorithm from
@@
  • Updating $\bm{b}^{*}$: Similar to $\bm{a}^{*}$, sample $\bm{b}^{*}$ jointly via Metropolis-Hastings algorithm from $f(\bm{b}^{*}|\bm{Y}_{1},\ldots,\bm{Y}_{T},\bm{a}^{*},\bm{c},\rho,\mu_{b},\sigma_{b}^{2},\rho_{b})$.
  • Updating $\bm{c}$: Similar to $\bm{a}^{*}$, and $\bm{b}^{*}$, sample $\bm{c}$ jointly via Metropolis-Hastings algorithm from $f(\bm{c}|\bm{Y}_{1},\ldots,\bm{Y}_{T},\bm{a}^{*},\bm{b}^{*},\rho,\mu_{c},\sigma_{c}^{2},\rho_{c})$ .
  • Updating $\rho$: Sample $\rho$ via the Metropolis-Hastings algorithm from
@@
  • Updating $\rho_{a}$: Sample $\rho_{a}$ via the Metropolis-Hastings algorithm from
@@
  • Updating $\rho_{b}$: Similarly, sample $\rho_{b}$ from $f(\rho_{b}|\bm{b}^{*},\mu_{b},\sigma_{b}^{2})$ via Metropolis-Hastings algorithm as $\rho_{a}$.
  • Updating $\rho_{c}$: Similarly, sample $\rho_{b}$ from $f(\rho_{c}|\bm{c},\mu_{c},\sigma_{c}^{2})$ via Metropolis-Hastings algorithm as $\rho_{a}$ and $\rho_{b}$.
  • Updating $Y^{*}_{it}s$: For a time point $t$, if $\mathcal{I}^{(t)}$ denotes the region indexes where the data are missing, we construct $\bm{Y}_{t}^{*}$ (vector of missing values) based on $i\in\mathcal{I}^{(t)}$ and the vector for the rest of the indexes as $\widetilde{\bm{Y}}_{t}$, i.e., based on all $i\in\bar{\mathcal{I}}^{(t)}=\{1,2,\dots,n\}\setminus\mathcal{I}^{(t)}$. Further, we transform all the components to $\textrm{Normal}(0,1)$ scale by the $\Phi^{-1}(F(Y_{it};a_{i},b_{i},c_{i}))$ transformation. Given that the transformed data follow a multivariate normal distribution, one can simulate the observations corresponding to the missing locations using the standard result of the conditional distribution for multivariate normal vectors, and then transform them back to the original scale using the Gaussian CDF and the gamma quantile function. However, as obtaining the precision matrix is easier in our case, we simulate using the result of the conditional distribution for multivariate normal vectors based on the precision matrix, which allows for faster analysis [rue2005gaussian]. Here, we use the partition of the precision matrix instead of inverting the parts of the variance-covariance matrix. Let the precision matrix $\bm{Q}$ of $\bm{Y}_{t}=({\bm{Y}_{t}^{*}}^{\intercal},\widetilde{\bm{Y}}_{t}^{\intercal})^{\intercal}$ be partitioned according to the missing observations as
    $\bm{Q}=\begin{bmatrix}\bm{Q}_{{\mathcal{I}}^{(t)},{\mathcal{I}}^{(t)}}&\bm{Q}_{{\mathcal{I}}^{(t)},\bar{\mathcal{I}}^{(t)}}\\ \\ \bm{Q}_{\bar{\mathcal{I}}^{(t)},{\mathcal{I}}^{(t)}}&\bm{Q}_{\bar{\mathcal{I}}^{(t)},\bar{\mathcal{I}}^{(t)}}\end{bmatrix}$.
    Since the means of the marginal distributions are zero, the conditional distribution of the missing data is $\bm{Y}_{t}^{*}\mid\widetilde{\bm{Y}}_{t}\sim\mathrm{Normal}_{|{\mathcal{I}}^{(t)}|}\left(-\bm{Q}_{{\mathcal{I}}^{(t)},{\mathcal{I}}^{(t)}}^{-1}\bm{Q}_{{\mathcal{I}}^{(t)},\bar{\mathcal{I}}^{(t)}}\widetilde{\bm{Y}}_{t},\ \bm{Q}_{{\mathcal{I}}^{(t)},{\mathcal{I}}^{(t)}}\right)$. Here $|{\mathcal{I}}^{(t)}|$ is the number of indices in the index set ${\mathcal{I}}^{(t)}$, i.e., the total number of missing observations at time point $t$.
For the univariate parameters updated via Metropolis–Hastings, alternative samplers such as the slice sampler with stepping-out [neal2003slice] could also be employed, which would reduce the need for proposal tuning; we retain the Metropolis-Hastings steps here due to its superior popularity in the environmental statistics literature.